Given that y'=x+y with y(0) = 1 find y(0.1).h=0.05 by Euler's modified method.


1
Expert's answer
2022-01-05T04:17:33-0500

Solution:

Given "y^{\prime}=x+y, y(0)=1, h=0.05, y(0.1)=?"

Here, "x_{0}=0, y_{0}=1, h=0.05, x_{n}=0.1"

"\begin{aligned}

&y^{\prime}=x+y \\

&\therefore f(x, y)=x+y

\end{aligned}"

Modified Euler method

"\begin{aligned}

&y_{m+1}=y_{m}+h f\left(x_{m}+\frac{1}{2} h, y_{m}+\frac{1}{2} h f\left(x_{m}, y_{m}\right)\right) \\

&f\left(x_{0}, y_{0}\right)=f(0,1)=1 \\

&\begin{array}{l}

1 \\

x_{0}+\frac{1}{2} h=0+\frac{0.05}{2}=0.025 \\

1 \\

y_{0}+\frac{1}{2} h f\left(x_{0}, y_{0}\right)=1+\frac{0.05}{2} \cdot 1=1.025 \\

f\left(x_{0}+\frac{1}{2} h, y_{0}+\frac{1}{2} h f\left(x_{0}, y_{0}\right)=f(0.025,1.025)=1.05\right. \\

\therefore y(0.05)=1.0525

\end{array}

\end{aligned}"

Again taking "\left(x_{1}, y_{1}\right)" in place of "\left(x_{0}, y_{0}\right)" and repeat the process

"\begin{aligned}

&f\left(x_{1}, y_{1}\right)=f(0.05,1.0525)=1.1025 \\

&x_{1}+\frac{1}{2} h=0.05+\frac{0.05}{2}=0.075 \\

&y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)=1.0525+\frac{0.05}{2} \cdot 1.1025=1.0801 \\

&f\left(x_{1}+\frac{1}{2} h, y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)=f(0.075,1.0801)=1.1551\right. \\

&y_{2}=y_{1}+h f\left(x_{1}+\frac{1}{2} h, y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)\right)=1.0525+0.05 \cdot 1.1551=1.1103 \\

&\therefore y(0.1)=1.1103

\end{aligned}"